OpenStudy (anonymous):

What is the value of log 5(subscript on 5) 19 rounded to the nearest hundredth?

OpenStudy (anonymous):

use the change of base forula: \[log_5(19)=\frac{\ln(19)}{ln(5)}\]

OpenStudy (anonymous):

or you can use \[log_5(19)=\frac{\log(19)}{log(5)}\]

OpenStudy (anonymous):

then use a calculator

OpenStudy (anonymous):

i get 2.469 rounded

OpenStudy (smurfy14):

1.68

OpenStudy (anonymous):

really? this time i got 1.83 rounded

OpenStudy (smurfy14):

oh ya 1.83 is right i thought the 19 was a 15 when i did it the first time